102) 2.37 moles of an ideal monatomic gas initially at 255 K undergoes this cycle: It is (1) heated at constant pressure to 655 K, (2) then allowed to cool at constant volume until its temperature returns to its initial value, (3) then compressed isothermally to its initial state. Find:

a. the net energy transferred as heat to the gas (excluding the energy transferred as heat out of the gas).

b. the net work done by the gas for the cycle.

c. the efficiency of this cycle.

d. What would be the efficiency of a Carnot Engine operating between the two heat reservoirs at 275 K and 645 K.
Answer #1

?, 25к вск 3 2 ск 5 TV, 2, 41v, и a) ASB beliabatic. isobaric ДО - hСо от - 2.37x Sx8 -3147 Чоо ДФ, 19769-27 W= nRAT - 2,37x

Denet - 19704-2-11822-51 : -4740 Adner ? 3141-66 37 b) W- 7881-67 +0-4740 lwa 3141-66 I] . c) dina 19764.2 ] n=w = 3.141-66